# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v2

@@ 2.1.1 Convergence analysis of ChFSI
Let the matrices $\textbf{E}\in\mathbb{C}^{m\times n}$ and $\textbf{E}_{\perp}\in\mathbb{C}^{m\times(m-n)}$ be given by $\textbf{E}=\begin{bmatrix}\textbf{I}_{n}\\ 0\end{bmatrix}$ and $\textbf{E}_{\perp}=\begin{bmatrix}0\\ \textbf{I}_{m-n}\end{bmatrix}$ where $\textbf{I}_{n}$ and $\textbf{I}_{m-n}$ are the $n\times n$ and $(m-n)\times(m-n)$ identity matrices respectively. Let $=\mathcal{R}(\textbf{E})$ E ) and $\mathcal{E}_{\perp}=\mathcal{R}(\textbf{E}_{\perp})$ where $\mathcal{R}(.)$ denotes range of a matrix. Then for any matrix $\textbf{X}\in\mathbb{C}^{m\times n}$ with $=\mathcal{R}(\textbf{X})$ X ) satisfying $\mathcal{E}_{\perp}\cap\mathcal{X}=\left\{0\right\}$ or equivalently $\text{rank}\left(\textbf{E}^{\dagger}\textbf{X}\right)=n$, we have
@@
Proof.
For a proof this statement we refer the reader to Zhu and Knyazev [67], Watkins [68]. ∎
For the sake of convenience in the subsequent convergence analysis, we define the Hermitian matrix $\hat{\textbf{H}}=\textbf{B}^{-\frac{1}{2}}\textbf{A}\textbf{B}^{-\frac{1}{2}}$ allowing us to rewrite the generalized eigenproblem $\textbf{A}\textbf{U}=\textbf{B}\textbf{U}\bm{\Lambda}$, as a Hermitian standard eigenvalue problem $\hat{\textbf{H}}\hat{\textbf{U}}=\hat{\textbf{U}}\bm{\Lambda}$, where $=\textbf{B}^{\frac{1}{2}}\textbf{U}$ U denotes the $m\times m$ unitary matrix with columns as eigenvectors of $\hat{\textbf{H}}$ and $\bm{\Lambda}$ denotes the diagonal matrix comprising eigenvalues of $\hat{\textbf{H}}$. Consider the partitioning of $\hat{\textbf{U}}$ as $\begin{bmatrix}\hat{\textbf{U}}_{1}&\hat{\textbf{U}}_{2}\end{bmatrix}$ where $\hat{\textbf{U}}_{1}$ is the $m\times n$ matrix whose columns are the eigenvectors corresponding to the lowest $n$ eigenvalues and $\hat{\textbf{U}}_{2}$ is the $m\times(m-n)$ matrix whose columns are the rest of the eigenvectors. We define the wanted eigenspace corresponding to the $n$ smallest eigenvectors of $\hat{\textbf{H}}$ as $\mathcal{S}=\mathcal{R}(\hat{\textbf{U}}_{1})$ and the unwanted eigenspace as $\mathcal{S}_{\perp}=\mathcal{R}(\hat{\textbf{U}}_{2})$. The trial subspace corresponding to $\hat{\textbf{H}}$ at the beginning of $i^{th}$ iteration is denoted as $\mathcal{S}^{(i)}=\mathcal{R}(\hat{\textbf{X}}^{(i)})$ where $\hat{\textbf{X}}^{(i)}=\textbf{B}^{\frac{1}{2}}\textbf{X}^{(i)}$. Further, the filtered subspace obtained at the end of this $i^{th}$ iteration is denoted as $\mathcal{S}^{(i+1)}=\mathcal{R}({\hat{\textbf{Y}}_{p}}^{(i)})$ where $\hat{\textbf{Y}}^{(i)}_{p}=\textbf{B}^{\frac{1}{2}}\textbf{Y}^{(i)}_{p}$. We note that using the relation $\textbf{B}^{\frac{1}{2}}C_{p}(\textbf{H})\textbf{B}^{-\frac{1}{2}}=C_{p}(\hat{\textbf{H}})$ and $\textbf{Y}_{p}^{(i)}=C_{p}(\textbf{H})\textbf{X}^{(i)}$ from Algorithm 1, we can conclude that $\hat{\textbf{Y}}^{(i)}_{p}=C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}$, i.e., $\mathcal{S}^{(i+1)}=C_{p}(\hat{\textbf{H}})\mathcal{S}^{(i)}$.
@@
Proof.
Since the maximum principal angle given by Definition 2.1 is invariant under unitary transformations, the tangent of the angles between the subspaces satisfy both $\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}=\tan{\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)},\hat{\textbf{U}}^{\dagger}\mathcal{S})}$ and $\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}=\tan{\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)},\hat{\textbf{U}}^{\dagger}\mathcal{S})}$. Since the space defined by $\hat{\textbf{U}}^{\dagger}\mathcal{S}$ is the column space $\mathcal{E}$ of the matrix $\textbf{E}=\begin{bmatrix}\textbf{I}_{n}\\ 0\end{bmatrix}$, we have the above angles between subspaces satisfying $\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}=\tan{\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)},\mathcal{E})}$ and $\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}=\tan{\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)},\mathcal{E})}$. Consider a partitioning of the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$ as $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\begin{bmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{bmatrix}$ where $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible, and consequently, we can write
@@
Subsequently from Lemma 2.1 we have $\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}=\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,$. Further, partitioning the diagonal matrix of eigenvalues $\bm{\Lambda}$, we have $\bm{\Lambda}=\begin{bmatrix}\bm{\Lambda}_{1}&0\\ 0&\bm{\Lambda}_{2}\end{bmatrix}$ where $\bm{\Lambda}_{1}$ is the $n\times n$ diagonal matrix comprising of the wanted eigenvalues and $\bm{\Lambda}_{2}$ is the $(m-n)\times(m-n)$ diagonal matrix comprising of the unwanted eigenvalues. We can now write
$\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)}=\hat{\textbf{U}}^{\dagger}C_{p}(\hat{\textbf{H}})\hat{\textbf{U}}\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)}=C_{p}(\bm{\Lambda})\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)}=\mathcal{R}\left(\begin{bmatrix}C_{p}(\bm{\Lambda}_{1})&0\\ 0&C_{p}(\bm{\Lambda}_{2})\end{bmatrix}\begin{bmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{bmatrix}\right)=\mathcal{R}\left(\begin{bmatrix}C_{p}(\bm{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}\\ C_{p}(\bm{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}\end{bmatrix}\right)$
Consequently from Lemma 2.1 we have $\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}=\lVert C_{p}(\bm{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,$ allowing us to write
$\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}\leq\lVert C_{p}(\bm{\Lambda}_{2})\rVert\,\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,\lVert C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,=\left\lvert\frac{C_{p}(\lambda_{n+1})}{C_{p}(\lambda_{n})}\right\rvert\,\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
3 Approximation Methods for Chebyshev Filtered Subspace Construction
@@ 3.1 Convergence of ChFSI with inexact subspace construction
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $\leq\left(\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,+\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}}\right)\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ (4)
where $\hat{\bm{\Delta}}_{p}^{(i)}=\underline{C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}}-C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}$
Proof.
Partitioning the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$, we have $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\begin{bmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{bmatrix}$ where $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible. Consequently, we can write
$\hat{\textbf{U}}^{\dagger}\underline{\mathcal{S}}^{(i+1)}$ $=\mathcal{R}\left(\hat{\textbf{U}}^{\dagger}\underline{C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}}\right)=\mathcal{R}\left(\hat{\textbf{U}}^{\dagger}C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}+\hat{\textbf{U}}^{\dagger}\hat{\bm{\Delta}}_{p}^{(i)}\right)=\mathcal{R}\left(C_{p}(\bm{\Lambda})\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}+\hat{\textbf{U}}^{\dagger}\hat{\bm{\Delta}}_{p}^{(i)}\right)$
$=\mathcal{R}\left(\begin{bmatrix}C_{p}(\bm{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}+\hat{\textbf{U}}_{1}^{\dagger}\hat{\bm{\Delta}}_{p}^{(i)}\\ C_{p}(\bm{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}+\hat{\textbf{U}}_{2}^{\dagger}\hat{\bm{\Delta}}_{p}^{(i)}\end{bmatrix}\right)$
@@
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $=\left\lVert\left(C_{p}(\bm{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}+\hat{\textbf{U}}_{2}^{\dagger}\hat{\bm{\Delta}}_{p}^{(i)}\right)\left(C_{p}(\bm{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}+\hat{\textbf{U}}_{1}^{\dagger}\hat{\bm{\Delta}}_{p}^{(i)}\right)^{-1}\right\rVert\,$
$=\left\lVert\left(C_{p}(\bm{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\bm{\Lambda}_{1})^{-1}+\hat{\textbf{U}}_{2}^{\dagger}\hat{\bm{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\bm{\Lambda}_{1})^{-1}\right)\left(\textbf{I}+\hat{\textbf{U}}_{1}^{\dagger}\hat{\bm{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\bm{\Lambda}_{1})^{-1}\right)^{-1}\right\rVert\,$
$\leq\frac{\lVert C_{p}(\bm{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,+\lVert\hat{\bm{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,}{1-\lVert\hat{\bm{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,}$ (5)
Note that in the last step, we have assumed that $\lVert\hat{\bm{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,<1$, and using the fact that $\lVert{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,=\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}$, a sufficient condition for this to be true is $\lvert C_{p}(\lambda_{n})\rvert\,\cos{\angle(\mathcal{S}^{(i)},\mathcal{S})}>\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,$. Upon further simplification using sub-multiplicative property of matrix spectral norms and triangle inequality, the inequality in Eq. 5 can be written as
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $\leq\frac{\lVert C_{p}(\bm{\Lambda}_{2})\rVert\,\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,\lVert C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,+\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\lVert{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,\lVert C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,}{1-\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\lVert{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,\lVert C_{p}(\bm{\Lambda}_{1})^{-1}\rVert\,}$
$\leq\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,}{\lvert C_{p}(\lambda_{n})\rvert\,\cos{\angle(\mathcal{S}^{(i)},\mathcal{S})}-\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,}$
$=\left(\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,+\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}}\right)\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
which proves the desired inequality in the theorem. ∎

### v3

@@ 2.1.1 Convergence analysis of ChFSI
Let the matrices $\textbf{E}\in\mathbb{C}^{m\times n}$ and $\textbf{E}_{\perp}\in\mathbb{C}^{m\times(m-n)}$ be given by $\textbf{E}=\Bigl{[}\begin{smallmatrix}\textbf{I}_{n}\\ 0\end{smallmatrix}\Bigr{]}$ and $\textbf{E}_{\perp}=\Bigl{[}\begin{smallmatrix}0\\ \textbf{I}_{m-n}\end{smallmatrix}\Bigl{]}$ where $\textbf{I}_{n}$ and $\textbf{I}_{m-n}$ are the $n\times n$ and $(m-n)\times(m-n)$ identity matrices respectively. Let $=\mathcal{R}(\textbf{E})$ E ) and $\mathcal{E}_{\perp}=\mathcal{R}(\textbf{E}_{\perp})$ where $\mathcal{R}(.)$ denotes range of a matrix. Then for any matrix $\textbf{X}\in\mathbb{C}^{m\times n}$ with $=\mathcal{R}(\textbf{X})$ X ) satisfying $\mathcal{E}_{\perp}\cap\mathcal{X}=\left\{0\right\}$ or equivalently $\text{rank}\left(\textbf{E}^{\dagger}\textbf{X}\right)=n$, we have
@@
Proof 2.3.
For proof of this statement, we refer the reader to [49, 41].
For the sake of convenience in the subsequent convergence analysis, we define the Hermitian matrix $\hat{\textbf{H}}=\textbf{B}^{-\frac{1}{2}}\textbf{A}\textbf{B}^{-\frac{1}{2}}$ allowing us to rewrite the generalized eigenproblem $\textbf{A}\textbf{U}=\textbf{B}\textbf{U}\boldsymbol{\Lambda}$, as a Hermitian standard eigenvalue problem $\hat{\textbf{H}}\hat{\textbf{U}}=\hat{\textbf{U}}\boldsymbol{\Lambda}$, where $=\textbf{B}^{\frac{1}{2}}\textbf{U}$ U denotes the $m\times m$ unitary matrix with columns as eigenvectors of $\hat{\textbf{H}}$ and $\boldsymbol{\Lambda}$ denotes the diagonal matrix comprising eigenvalues of $\hat{\textbf{H}}$. Consider the partitioning of $\hat{\textbf{U}}$ as $\begin{bmatrix}\hat{\textbf{U}}_{1}&\hat{\textbf{U}}_{2}\end{bmatrix}$ where $\hat{\textbf{U}}_{1}$ is the $m\times n$ matrix whose columns are the eigenvectors corresponding to the lowest $n$ eigenvalues and $\hat{\textbf{U}}_{2}$ is the $m\times(m-n)$ matrix whose columns are the rest of the eigenvectors. We define the wanted eigenspace corresponding to the $n$ smallest eigenvectors of $\hat{\textbf{H}}$ as $\mathcal{S}=\mathcal{R}(\hat{\textbf{U}}_{1})$ and the unwanted eigenspace as $\mathcal{S}_{\perp}=\mathcal{R}(\hat{\textbf{U}}_{2})$. The trial subspace corresponding to $\hat{\textbf{H}}$ at the beginning of $i^{th}$ iteration is denoted as $\mathcal{S}^{(i)}=\mathcal{R}(\hat{\textbf{X}}^{(i)})$ where $\hat{\textbf{X}}^{(i)}=\textbf{B}^{\frac{1}{2}}\textbf{X}^{(i)}$. Further, the filtered subspace obtained at the end of the $i^{th}$ iteration is denoted as $\mathcal{S}^{(i+1)}=\mathcal{R}({\hat{\textbf{Y}}_{p}}^{(i)})$ where $\hat{\textbf{Y}}^{(i)}_{p}=\textbf{B}^{\frac{1}{2}}\textbf{Y}^{(i)}_{p}$. We note that using the relation $\textbf{B}^{\frac{1}{2}}C_{p}(\textbf{H})\textbf{B}^{-\frac{1}{2}}=C_{p}(\hat{\textbf{H}})$ and $\textbf{Y}_{p}^{(i)}=C_{p}(\textbf{H})\textbf{X}^{(i)}$ from Algorithm 1, we can conclude that $\hat{\textbf{Y}}^{(i)}_{p}=C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}$, i.e., $\mathcal{S}^{(i+1)}=C_{p}(\hat{\textbf{H}})\mathcal{S}^{(i)}$.
@@
Proof 2.5.
Since the maximum principal angle given by Definition 2.1 is invariant under unitary transformations, we have ${\angle(\mathcal{S}^{(i+1)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)},\hat{\textbf{U}}^{\dagger}\mathcal{S})}$ and also ${\angle(\mathcal{S}^{(i)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)},\hat{\textbf{U}}^{\dagger}\mathcal{S})}$. Since the space defined by $\hat{\textbf{U}}^{\dagger}\mathcal{S}$ is the same as $=\mathcal{R}(\textbf{E})$ E ), we have, ${\angle(\mathcal{S}^{(i+1)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)},\mathcal{E})}$ and ${\angle(\mathcal{S}^{(i)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)},\mathcal{E})}$. Consider a partitioning of the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$ as $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\Bigl{[}\begin{smallmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{smallmatrix}\Bigr{]}$ where $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible, and consequently, we can write
@@
Subsequently from Lemma 2.2 we have $\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}=\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,$. Further, partitioning the diagonal matrix of eigenvalues $\boldsymbol{\Lambda}$, we have $\boldsymbol{\Lambda}=\Bigl{[}\begin{smallmatrix}\boldsymbol{\Lambda}_{1}&0\\ 0&\boldsymbol{\Lambda}_{2}\end{smallmatrix}\Bigr{]}$ where $\boldsymbol{\Lambda}_{1}$ is the $n\times n$ diagonal matrix comprising of the wanted eigenvalues and $\boldsymbol{\Lambda}_{2}$ is the $(m-n)\times(m-n)$ diagonal matrix comprising of the unwanted eigenvalues. We can now write
$\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)}=\hat{\textbf{U}}^{\dagger}C_{p}(\hat{\textbf{H}})\mathcal{S}^{(i)}=C_{p}(\boldsymbol{\Lambda})\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)}\\ =\mathcal{R}\left(\begin{bmatrix}C_{p}(\boldsymbol{\Lambda}_{1})&0\\ 0&C_{p}(\boldsymbol{\Lambda}_{2})\end{bmatrix}\begin{bmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{bmatrix}\right)=\mathcal{R}\left(\begin{bmatrix}C_{p}(\boldsymbol{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}\\ C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}\end{bmatrix}\right)$
From Lemma 2.2 we now have $\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}=\lVert C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,$, allowing us to write
$\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}\leq\lVert C_{p}(\boldsymbol{\Lambda}_{2})\rVert\,\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,\lVert C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,=\left\lvert\frac{C_{p}(\lambda_{n+1})}{C_{p}(\lambda_{n})}\right\rvert\,\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
3 Approximate Chebyshev Filtered Subspace Construction
@@ 3.1 Convergence of ChFSI with inexact subspace construction
(3) $\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $\leq\left(\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}}\right)\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
where $\hat{\boldsymbol{\Delta}}_{p}^{(i)}=\underline{C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}}-C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}$
Proof 3.2.
Partitioning the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$, we have $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\Bigl{[}\begin{smallmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{smallmatrix}\Bigr{]}$ where $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible. Consequently, we can write
$\hat{\textbf{U}}^{\dagger}\underline{\mathcal{S}}^{(i+1)}$ $=\mathcal{R}\left(\hat{\textbf{U}}^{\dagger}\underline{C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}}\right)=\mathcal{R}\left(\hat{\textbf{U}}^{\dagger}C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}+\hat{\textbf{U}}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)$
$=\mathcal{R}\left(C_{p}(\boldsymbol{\Lambda})\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}+\hat{\textbf{U}}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)=\mathcal{R}\left(\begin{bmatrix}C_{p}(\boldsymbol{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}+\hat{\textbf{U}}_{1}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\\ C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}+\hat{\textbf{U}}_{2}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\end{bmatrix}\right)$
@@
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $=\left\lVert\left(C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}+\hat{\textbf{U}}_{2}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)\left(C_{p}(\boldsymbol{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}+\hat{\textbf{U}}_{1}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)^{-1}\right\rVert\,$
(4) $\leq\frac{\lVert C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,}{1-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,}$
Note that in the last step, we have assumed that $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,<1$, and using the fact that $\lVert{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,=\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}$, a sufficient condition for this to be true is $\lvert C_{p}(\lambda_{n})\rvert\,\cos{\angle(\mathcal{S}^{(i)},\mathcal{S})}>\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$. Upon further simplification using sub-multiplicative property of matrix spectral norms and triangle inequality, the inequality in Eq. 4 can be written as
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $\leq\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,}{\lvert C_{p}(\lambda_{n})\rvert\,\cos{\angle(\mathcal{S}^{(i)},\mathcal{S})}-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,}$
$=\left(\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}}\right)\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
which proves the desired inequality in the theorem.